Add a term to make the expression a perfect square. Then factor the perfect square.
x^2 + 1/7x

Answer by Boreal(15235)

One adds the square of half the x value, which would be (1/14)^2=1/196
(x+1/14)^2=x^2+(2/14) or (1/7)x + 1/196